OpenStudy (anonymous):

Lecture #5 textbook homework check. A&E Chapter 6 problems 76 and 78 do not have answers at the back of the book. Neither of these questions is difficult, but for completeness I would like to swap answers.

OpenStudy (anonymous):

A&E 6-76) a) 1s has 2 electrons b) 4p has 6 electrons c) 5d has 10 electrons d) 4f has 14 electrons A&E 6-78) a) n=1, l=0, m=0, s=+1/2 is 1s b) n=2, l=1, m=0, s=+1/2 is 2p c) n=3, l=2, m=0, s=+1/2 is 3d d) n=4, l=3, m=3, s=+1/2 is 4f